OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. x^2 + 8x - 5 = 0

OpenStudy (anonymous):

@phi

OpenStudy (phi):

do you know the quadratic formula?

OpenStudy (anonymous):

0.58257569495584, -8.58257569495584.

OpenStudy (anonymous):

ya \[-b \pm \sqrt{b ^{2}-4ac}\]

OpenStudy (anonymous):

i gave u the answer

OpenStudy (anonymous):

ya but I need to know how to do it

OpenStudy (phi):

If you are fuzzy on using it, watch this when you have time http://www.khanacademy.org/math/algebra/quadratics/quadratic_formula/v/using-the-quadratic-formula meanwhile, you want to match the letters a, b and c to your problem

OpenStudy (anonymous):

I KNOW

OpenStudy (phi):

x^2 + 8x - 5 = 0 you "pattern match" that to a x^2 + b x + c=0 what do you get for a, b and c ?

OpenStudy (phi):

btw, we want to divide by 2a in the formula \[ (-b \pm \sqrt{b ^{2}-4ac})/2a \]

OpenStudy (phi):

First step is give the numbers for a,b and c can you do that ?

OpenStudy (anonymous):

A=1 B=8 C=-5

OpenStudy (anonymous):

Am I right

OpenStudy (anonymous):

@ProfessorMath can u help me

OpenStudy (anonymous):

@Callisto will u hel me

OpenStudy (callisto):

a=1, b=8, c=-5 Correct. Now substitute these values into the quadratic formula \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]to find x.

OpenStudy (anonymous):

\[-8\pm \sqrt{8^{2}-4*1*-5}\]

OpenStudy (callisto):

No.. You need the denominator (2a) too!

OpenStudy (anonymous):

\[-8\pm \sqrt{84-4*1*5}\div2*1\]

OpenStudy (anonymous):

\[-8\pm \sqrt{80*1*5\div2}\]

OpenStudy (callisto):

8^2 = 64 \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}=\frac{-8\pm \sqrt{8^2-4(1)(-5)}}{2(1)}=\frac{-8\pm \sqrt{64-4(1)(-5)}}{2(1)}=...?\]

OpenStudy (anonymous):

ya I get stuck after this

OpenStudy (callisto):

Simplify the value under the square root first. What is 64 - 4(1)(-5)?

OpenStudy (anonymous):

84

OpenStudy (anonymous):

Can one of yal help me

OpenStudy (callisto):

Yes. So, now, we get \[x=\frac{-8\pm \sqrt{84}}{2(1)}\] 84 = (2^2)x3x7, so \[x=\frac{-8\pm \sqrt{2^2\times 3\times 7}}{2(1)}=\frac{-8\pm 2\sqrt{3\times 7}}{2(1)}\]Can you simplify it?

OpenStudy (anonymous):

they both equal 84

OpenStudy (callisto):

Hmm... they =?

OpenStudy (anonymous):

idk im getting confused

OpenStudy (callisto):

Confused at?

OpenStudy (anonymous):

I have to find what X=

OpenStudy (callisto):

That's what we have been working on.

OpenStudy (anonymous):

o wow this takes a long time

OpenStudy (anonymous):

so I go t \[x=-8\pm \sqrt{84} \div2\]

OpenStudy (callisto):

\[x=(-8\pm \sqrt{84})\div2\]

OpenStudy (anonymous):

ya wats next

OpenStudy (callisto):

Simplify it.

OpenStudy (anonymous):

okay so its \[-8\pm9.16515\div2\]

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

\[-8\pm4.58258\] this is after I div 9.16515 by 2

OpenStudy (callisto):

\[(-8\pm 9.16515)\div2\]Add/subtract before you divide.

OpenStudy (anonymous):

the \[-8+9.16515 \div2=0.58258\] that's add the \[-8-9.16515\div2=-8.58258\] that's subtract

OpenStudy (anonymous):

Is that right @Callisto

OpenStudy (callisto):

Yes. But remember to put brackets for the addition/subtraction part.

OpenStudy (anonymous):

what would x=